OpenStudy (anonymous):

THIS IS REALLY EASYYY. but Idk the answer. how do you find x in this: 2/x-1 - 3/x

OpenStudy (anonymous):

break into like terms 2/x - 3/x = 1 then simply minus the fractions

OpenStudy (anonymous):

are you sureee? I mean there's like an x-1 there..

OpenStudy (anonymous):

or not.

OpenStudy (anonymous):

2/x-1-3/x=0 =>2x-3(x-1)/x(x-1)=0 2x-3x+3/x(x-1)=0 -x+3/x^2-x=0 =>3-x/x^2-x=0 =>3-x=0 =>3=x x=3

OpenStudy (anonymous):

i agree x = 3.

OpenStudy (anonymous):

Thanks a lot but could you explain from =>3-x/x^2-x=0 onwards?

OpenStudy (anonymous):

see if u can see what i have attached.